The risk-free rate is 3.67% and the market risk premium is 6.46%. A stock with a β of 1.64 just paid a dividend of $1.74. The dividend is expected to grow at 20.87% for three years and then grow at 4.52% forever. What is the value of the stock?

Please round to 2 decimal places.
Answer #1

required return=risk free rate+beta*market risk premium=3.67%+1.64*6.46%=14.26%

Value of the stock=Present value of dividends

=1.74*((1+g)/(1+r))+1.74*((1+g)/(1+r))^2+1.74*((1+g)/(1+r))^3+1.74*((1+g)/(1+r))^3*(1+t)/(r-g)

=1.74*((1+20.87%)/(1+14.26%))+1.74*((1+20.87%)/(1+14.26%))^2+1.74*((1+20.87%)/(1+14.26%))^3+1.74*((1+20.87%)/(1+14.26%))^3*(1+4.52%)/(14.26%-4.52%)

=27.93817651